News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 2 April 2004 8.30am EST

During the last week Chandra completed the observing schedule as planned.

The IEPHIN instrument exhibited anomalous behavior on Mar 26 during a period of elevated operating temperature. The power supply voltage dropped as it became current limited due to the high temperature and returned to its nominal value once the unit cooled following a maneuver to a new attitude. This behavior has been seen before and mission planning constraints have been in use to avoid high temperatures. In this case the actual temperature exceeded the projected value by a small margin, and as a result the constraints will be tightened accordingly.

A real-time procedure was completed on Mar 29 to perform a manual unload of momentum. The unload was performed at an attitude that ensured a cool starting temperature and thruster performance was nominal throughout the unload.

The schedule of targets for the next week is shown below and includes a series of calibration observations of the oxygen rich supernova remnant E0102-72, and an observation of the Earth.

------------------------------------------
RX J1914.4+2456        ACIS-S       Apr 4
E0102-72               ACIS-S       Apr 5
ESO198-G24             ACIS-S/HETG
Radiation Belts
GRS 1915+105           ACIS-S/HETG  Apr 6
SN2004C                ACIS-S       Apr 7
F1215-1639             ACIS-S
F1350-4593             ACIS-S
ESO198-G24             ACIS-S/HETG
Radiation Belts                     Apr 8
E0102-72 (4 obs)       ACIS-S       Apr 9
NGC 4395               ACIS-S
Earth                  HRC-I        Apr 10
E0102-72 (2 obs)       ACIS-S
E0102-72               ACIS-I
NGC 4395               ACIS-S

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports